the sum of two numbers is 10. Three times the lesser number is 5 less that twice the greater number What are the numbers?
Answer by checkley77(12844)

X+Y=10 OR X=10-Y
3X=2Y-5
3(10-Y)=2Y-5
30-3Y=2Y-5
-3Y-2Y=-5-30
-5Y=-35
Y=-25/-5
Y=7 ANS. FOR THE LARGER NUMBER.
X=10-7=3 ANS. FOR THE LESSER NUMBER.
PROOF:
3*3=2*7-5
9=14-5
9=9